# What Factors Affect LLMs and RLLMs in Financial Question Answering?

### v3

@@ Abstract
Recently, the development of large language models (LLMs) and reasoning large language models (RLLMs) have gained considerable attention from many researchers. RLLMs enhance the reasoning capabilities of LLMs through Long Chain-of-Thought (Long CoT) processes, significantly improving the performance of LLMs in addressing complex problems. However, there are few works that systematically explore what methods can fully unlock the performance of LLMs and RLLMs within the financial domain. To investigate the impact of various methods on LLMs and RLLMs, we utilize five LLMs and three RLLMs to assess the effects of prompting methods, agentic frameworks, and multilingual alignment methods on financial question-answering tasks. Our research findings indicate: (1) Current prompting methods and agent frameworks enhance the performance of LLMs in financial question answering by simulating Long CoT; (2) RLLMs possess inherent Long CoT capabilities, which limits the effectiveness of conventional methods in further enhancing their performance; (3) Current advanced multilingual alignment methods primarily improve the multilingual performance of LLMs by extending the reasoning length, which yields minimal benefits for RLLMs. Additionally, we discuss strategies for enhancing the performance of LLMs and RLLMs in financial question answering, which may serve as a inspiration for future improvements. We hope that this study can serve as an important reference for LLMs and RLLMs in the field of financial question answering.
@@ 1 Introduction
Recently, large language models (LLMs) have significantly advanced the field of natural language processing (NLP), and more and more researchers utilize LLMs to solve complex tasks in various domains (Zhou et al., 2023; Nie et al., 2024; Qin et al., 2024; Chang et al., 2024; Wang et al., 2025a). Furthermore, some researchers propose prompting-based and agentic frameworks to enhance the capabilities of LLMs (Wei et al., 2022; Yao et al., 2023a; Zhang et al., 2024b; Yao et al., 2023b; Li et al., 2023; Hong et al., 2024), which further expands the capabilities of LLMs in tasks in different fields. Particularly, to enhance the reasoning capabilities of LLMs in addressing complex problems, researchers introduce reasoning large language models (RLLMs). The application of the Long Chain-of-Thought (Long CoT) fully unlocks the understanding ability of RLLMs (Chen et al., 2025). Long CoT entails deeper reasoning, reflective analysis, and a more extensive exploration of logical structures (Gandhi et al., 2025), which enables RLLMs to perform on par with experts in certain domains.
Researchers have explored the application of LLMs in financial question answering. Wu et al. (2023) present BloombergGPT, a 50 billion parameter LLM for financial domain. Fatemi & Hu (2024) propose a multi-agent framework to enhance the performance of LLMs in financial question answering. Xie et al. (2024) introduce a novel financial benchmark, FinBen, including 42 datasets spanning 24 financial tasks. Srivastava et al. (2024) evaluate LLMs in four financial tabular question-answering datasets. Xue et al. (2024) build a benchmark for financial multilingual multi-modal question answering named FAMMA, which is a challenging benchmark for multi-modal LLMs. Wang et al. (2025c) propose FinSage, which solves the financial filings question answering task through retrieval augmented generation (RAG).
Although LLMs have made significant progress in the field of financial question answering, existing work still mainly focuses on enhancing their performance within this domain. With the advancement of LLMs and RLLMs, the factors influencing their financial question-answering capabilities remain under-explored. Inspired by this gap, we systematically investigate the following issues: ($i$) the impact of prompting methods on LLMs/RLLMs; ($ii$) the influence of agentic frameworks on LLMs/RLLMs; and (iiiiii) whether multilingual alignment methods can enhance the multilingual financial question-answering abilities of LLMs/RLLMs. We utilize five LLMs and three RLLMs, conducting detailed experiments across seven representative methods.
@@ 3.1 Exploration of prompting methods
Current research extensively demonstrates that different prompting methods can significantly alter the performance of LLMs (Wei et al., 2022; Wang et al., 2023; Zhang et al., 2024b). The prompting method modifies certain aspects of the thinking prompt while ensuring the question remains intact, enabling LLMs to solve the question by following the structured steps of the thinking prompt.
We select three representative prompting methods to investigate the impact of varying prompt word changes on the performance of LLMs and RLLMs: Direct, Zero-shot CoT (Kojima et al., 2022), and Plan-and-Solve (Wang et al., 2023). These prompting methods can be defined as:
@@ 3.2 Exploration of agentic frameworks
To further enhance the performance of LLMs, researchers have started to incorporate agentic systems. Specifically, by enabling LLMs to assume different roles and engage in continuous communication and cooperation, their performance can be significantly improved (Luo et al., 2025a).
We select two agentic frameworks, and aim to explore the impact of LLMs and RLLMs in different agent collaborations: Self-Refine (Madaan et al., 2023) and S3 Agent (Wang et al., 2024). These agentic frameworks can be defined as:
@@ 3.3 Exploration of multilingual alignment methods
We evaluate the performance of four standard multilingual approaches, including state-of-the-art cross-lingual prompting methods, on multilingual alignment: Direct, Translate-en (Shi et al., 2023), and Cross-lingual Prompting (Qin et al., 2023).
@@
  • Translate-en: Translate-en employs translation to convert the local language question into English, thereby preventing performance degradation of LLMs due to insufficient multilingual capabilities.
@@
Method Arithmetic Non-Arithmetic Overall Overall Easy Medium Hard Overall Easy Medium Hard Llama-3.1-8B-Instruct (Grattafiori et al., 2024) Direct (Grattafiori et al., 2024) 10.88 12.66 9.89 9.28 23.44 22.45 32.10 20.27 16.50 Zero-shot CoT (Kojima et al., 2022) 13.20 16.15 10.95 11.07 31.72 32.62 34.73 29.95 21.49 Plan-and-Solve (Wang et al., 2023) 16.09 18.99 14.84 13.17 29.77 32.62 30.52 27.92 22.21 Self-Refine (Madaan et al., 2023) 17.95 22.70 13.07 15.56 32.06 34.74 35.26 29.27 24.26 S3 Agent (Wang et al., 2024) 19.62 22.27 15.54 19.46 30.80 37.28 32.63 26.57 24.62 GPT-4o-mini (Hurst et al., 2024) Direct (Hurst et al., 2024) 32.46 41.70 30.38 21.55 42.52 50.42 50.52 34.90 36.96 Zero-shot CoT (Kojima et al., 2022) 34.13 41.04 34.27 24.55 45.17 56.35 52.10 36.26 39.07 Plan-and-Solve (Wang et al., 2023) 36.18 44.97 33.92 26.04 45.74 56.77 51.05 37.61 40.46 Self-Refine (Madaan et al., 2023) 33.30 41.26 33.21 22.45 42.06 52.11 45.78 35.13 37.22 S3 Agent (Wang et al., 2024) 30.88 38.42 28.26 22.75 41.72 51.69 46.84 34.23 35.73 Gemini-1.5-flash (Team et al., 2024) Direct (Team et al., 2024) 37.20 47.16 35.33 25.14 45.28 58.05 49.47 36.71 40.82 Zero-shot CoT (Kojima et al., 2022) 37.39 46.72 36.04 25.74 46.09 58.89 51.57 36.93 41.28 Plan-and-Solve (Wang et al., 2023) 37.76 46.50 39.57 24.25 47.12 55.50 55.26 39.18 41.95 Self-Refine (Madaan et al., 2023) 34.41 42.35 33.92 23.95 42.52 52.96 50.00 33.78 38.04 S3 Agent (Wang et al., 2024) 37.20 45.85 34.62 27.54 44.59 57.20 50.52 35.36 40.51 Qwen-2.5-32B (Yang et al., 2024) Direct (Yang et al., 2024) 41.86 50.65 40.98 30.53 48.62 59.32 56.31 39.63 44.88 Zero-shot CoT (Kojima et al., 2022) 42.32 50.21 40.98 32.63 50.80 59.32 60.00 42.34 46.11 Plan-and-Solve (Wang et al., 2023) 39.44 44.97 43.46 28.44 49.77 60.59 55.78 41.44 44.06 Self-Refine (Madaan et al., 2023) 41.58 49.56 42.75 29.64 49.65 57.62 54.73 43.24 45.19 S3 Agent (Wang et al., 2024) 42.51 51.52 41.34 31.13 48.85 61.44 56.31 38.96 45.34 DeepSeek-V3 (Liu et al., 2024) Direct (Liu et al., 2024) 56.27 65.50 58.65 41.61 62.06 73.72 69.47 52.70 58.86 Zero-shot CoT (Kojima et al., 2022) 56.18 63.97 59.71 42.51 61.83 71.18 70.52 53.15 58.71 Plan-and-Solve (Wang et al., 2023) 55.16 63.97 56.18 42.21 63.33 70.76 71.05 56.08 58.81 Self-Refine (Madaan et al., 2023) 53.86 61.79 55.83 41.31 59.54 69.06 64.73 52.25 56.40 S3 Agent (Wang et al., 2024) 54.69 61.57 56.89 43.41 59.42 71.18 66.31 50.22 56.81 DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) Direct (Guo et al., 2025) 50.41 60.48 50.17 36.82 57.12 65.25 64.73 49.54 53.41 Zero-shot CoT (Kojima et al., 2022) 51.25 60.91 51.94 37.42 56.55 68.64 63.15 47.29 53.62 Plan-and-Solve (Wang et al., 2023) 51.16 60.91 50.53 38.32 55.86 66.52 62.63 47.29 53.26 Self-Refine (Madaan et al., 2023) 44.65 52.18 47.70 31.73 52.06 57.20 60.00 45.94 47.96 S3 Agent (Wang et al., 2024) 51.25 59.60 51.94 39.22 58.04 69.49 65.26 48.87 54.29 Qwen-3-14B (Yang et al., 2025) Direct (Yang et al., 2025) 54.88 66.59 58.65 35.62 59.08 72.03 70.52 47.29 56.76 Zero-shot CoT (Kojima et al., 2022) 55.72 68.34 56.18 38.02 59.19 69.91 70.52 48.64 57.27 Plan-and-Solve (Wang et al., 2023) 54.97 66.81 58.65 35.62 57.81 71.18 68.94 45.94 56.24 Self-Refine (Madaan et al., 2023) 54.41 65.93 56.89 36.52 58.27 71.61 66.84 47.52 56.14 S3 Agent (Wang et al., 2024) 54.13 65.06 57.24 36.52 58.50 71.18 67.36 47.97 56.09 O4-mini (OpenAI, 2025) Direct (OpenAI, 2025) 61.30 70.08 65.01 46.10 70.22 77.54 75.78 63.96 65.29 Zero-shot CoT (Kojima et al., 2022) 62.60 69.86 64.66 50.89 71.37 77.11 77.89 65.54 66.52 Plan-and-Solve (Wang et al., 2023) 61.67 69.86 64.31 48.20 72.41 77.96 77.89 67.11 66.47 Self-Refine (Madaan et al., 2023) 61.20 69.21 60.77 50.59 70.34 75.00 74.73 65.99 65.29 S3 Agent (Wang et al., 2024) 61.76 69.65 62.19 50.59 71.49 76.69 75.78 66.89 66.11
@@ 4.1 Experimental Settings
We use the standard financial question answering benchmark FAMMA (Xue et al., 2024) for experiments. FAMMA is a benchmark for financial multilingual multi-modal question answering, which includes English, Chinese, and French. Since current RLLMs are unimodal and cannot process multimodal information, we utilize the Basic Txt dataset in FAMMA, which converts multimodal data into textual information using OCR. The final dataset used for evaluation comprises 1,945 entries.
We conduct experiments on seven backbones, including five LLMs and three RLLMs: Meta-Llama-3.1-8B-Instruct (Grattafiori et al., 2024), GPT-4o-mini (Hurst et al., 2024), Gemini-1.5-flash (Team et al., 2024), Qwen-2.5-32B (Yang et al., 2024), DeepSeek-V3 (Liu et al., 2024), DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025), thing mode of Qwen-3-14B (Yang et al., 2025), O4-mini (OpenAI, 2025). Detailed settings can be found in the Appendix LABEL:appendix:details.
@@ 4.2 Analysis of Prompting Methods
Larger LLMs exhibit increased robustness to various prompting methods. It is evident that for larger LLMs, such as Gemini-1.5-flash, Qwen-2.5-32B, and DeepSeek-V3, the performance gap between different prompting methods is not significant, with performance fluctuations remaining under 2%. This indicates that LLMs with superior performance are less affected by input prompts, demonstrating greater robustness to variations in input.
Performance Direct Translate-en (Shi et al., 2023) Self-Refine (Madaan et al., 2023) S3 Agent (Wang et al., 2024) CLP (Qin et al., 2023) Meta-Llama-3.1-8B-Instruct (Grattafiori et al., 2024) Chinese 16.54 16.54 (+0.00) 18.97 (+2.43) 19.22 (+2.68) 23.35 (+6.81) French 15.38 16.66 (+1.28) 13.46 (-1.92) 21.15 (+5.77) 17.30 (+1.92) Overall (Chinese+French) 16.22 16.57 (+0.35) 17.46 (+1.24) 19.75 (+3.53) 21.69 (+5.47) GPT-4o-mini (Hurst et al., 2024) Chinese 33.09 34.54 (+1.45) 34.79 (+1.70) 31.63 (-1.46) 38.19 (+5.10) French 28.84 30.76 (+1.92) 28.84 (+0.00) 32.69 (+3.85) 32.05 (+3.21) Overall (Chinese+French) 31.92 33.50 (+1.58) 33.15 (+1.23) 31.92 (+0.00) 36.50 (+4.58) Gemini-1.5-flash (Team et al., 2024) Chinese 40.38 38.68 (-1.70) 36.73 (-3.65) 39.65 (-0.73) 37.22 (-3.16) French 27.56 29.48 (+1.92) 26.92 (-0.64) 27.56 (+0.00) 29.48 (+1.92) Overall (Chinese+French) 36.86 36.15 (-0.71) 34.03 (-2.83) 36.33 (-0.53) 35.09 (-1.77) Qwen-2.5-32B (Yang et al., 2024) Chinese 41.84 44.52 (+2.68) 44.28 (+2.44) 42.82 (+0.98) 43.30 (+1.46) French 29.48 29.48 (+0.00) 33.97 (+4.49) 35.89 (+6.41) 35.25 (+5.77) Overall (Chinese+French) 38.44 40.38 (+1.94) 41.44 (+3.00) 40.91 (+2.47) 41.09 (+2.65) DeepSeek-V3 (Liu et al., 2024) Chinese 57.42 58.63 (+1.21) 55.96 (-1.46) 57.66 (+0.24) 61.31 (+3.89) French 50.00 50.00 (+0.00) 47.43 (-2.57) 44.23 (-5.77) 57.69 (+7.69) Overall (Chinese+French) 55.37 56.26 (+0.89) 53.61 (-1.76) 53.96 (-1.41) 60.31 (+4.94) DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) Chinese 48.41 49.87 (+1.46) 46.47 (-1.94) 52.06 (+3.25) 50.85 (+2.44) French 49.35 42.94 (-6.41) 44.87 (-1.16) 49.35 (+0.00) 52.56 (+3.21) Overall (Chinese+French) 48.67 47.97 (-0.70) 46.03 (-2.64) 51.32 (+2.65) 51.32 (+2.65) Qwen-3-14B (Yang et al., 2025) Chinese 57.66 56.69 (-0.97) 55.23 (-2.43) 55.47 (-2.19) 53.77 (-3.89) French 48.71 50.00 (+1.29) 42.3 (-6.41) 48.71 (+0.00) 51.28 (+2.57) Overall (Chinese+French) 55.20 54.85 (-0.35) 51.67 (-3.53) 53.61 (-1.59) 53.08 (-2.12) O4-mini (OpenAI, 2025) Chinese 65.45 65.93 (+0.48) 64.23 (-1.22) 67.63 (+2.18) 64.54 (-0.91) French 62.17 61.53 (-0.64) 60.25 (-1.92) 61.53 (-0.64) 63.46 (+1.29) Overall (Chinese+French) 64.55 64.72 (+0.17) 63.13 (-1.42) 65.69 (+1.14) 64.24 (-0.31)
@@
For RLLMs, competitive results can be achieved without the use of prompting methods. Benefiting from the Long CoT process, RLLMs can generate complex reasoning independently, without deliberate guidance from prompting methods. This renders prompting, which is effective for LLMs, less effective for RLLMs and may even diminish their performance. As shown in Table 1, Plan-and-Solve demonstrates poor performance on the three RLLMs.

### v4

@@ Abstract
Recently, large language models (LLMs) and reasoning large language models (RLLMs) have gained considerable attention from many researchers. RLLMs enhance the reasoning capabilities of LLMs through Long Chain-of-Thought (Long CoT) processes, significantly improving the performance of LLMs in addressing complex problems. However, there are few works that systematically explore what methods can fully unlock the performance of LLMs and RLLMs within the financial domain. To investigate the impact of various methods on LLMs and RLLMs, we utilize five LLMs and four RLLMs to assess the effects of prompting methods, agentic frameworks, and multilingual alignment methods on financial question-answering tasks. Our research findings indicate: (1) Current prompting methods and agent frameworks enhance the performance of LLMs in financial question answering by simulating Long CoT; (2) RLLMs possess inherent Long CoT capabilities, which limits the effectiveness of conventional methods in further enhancing their performance; (3) Current advanced multilingual alignment methods primarily improve the multilingual performance of LLMs by extending the reasoning length, which yields minimal benefits for RLLMs. Additionally, we discuss strategies for enhancing the performance of LLMs and RLLMs in financial question answering, which may serve as a inspiration for future improvements. We hope that this study can serve as an important reference for LLMs and RLLMs in the field of financial question answering.
What Factors Affect LLMs and RLLMs in Financial Question Answering?
[…]
@@ 1 Introduction
Recently, large language models (LLMs) have significantly advanced the field of natural language processing (NLP), and more and more researchers utilize LLMs to solve complex tasks in various domains (zhou2023survey; nie2024survey; qin2024large; chang2024survey; peng2025large). Furthermore, some researchers propose prompting-based and agentic frameworks to enhance the capabilities of LLMs (wei2022chain; yao2023tree; zhang-etal-2024-wrong; yao2023react; li2023camel; hong2024metagpt), which further expands the capabilities of LLMs in tasks in different fields. Particularly, to enhance the reasoning capabilities of LLMs in addressing complex problems, researchers introduce reasoning large language models (RLLMs). The application of the Long Chain-of-Thought (Long CoT) fully unlocks the understanding ability of RLLMs (chen2025towards). Long CoT entails deeper reasoning, reflective analysis, and a more extensive exploration of logical structures (gandhi2025cognitive), which enables RLLMs to perform on par with experts in certain domains.
Researchers have explored the application of LLMs in financial question answering. wu2023bloomberggpt present BloombergGPT, a 50 billion parameter LLM for financial domain. 10.1145/3677052.3698686 propose a multi-agent framework to enhance the performance of LLMs in financial question answering. xie2024finben introduce a novel financial benchmark, FinBen, including 42 datasets spanning 24 financial tasks. srivastava2024evaluating evaluate LLMs in four financial tabular question-answering datasets. xue2024famma build a benchmark for financial multilingual multi-modal question answering named FAMMA, which is a challenging benchmark for multi-modal LLMs. wang2025finsage propose FinSage, which solves the financial filings question answering task through retrieval augmented generation (RAG).
Although LLMs have made significant progress in the field of financial question answering, existing work still mainly focuses on enhancing their performance within this domain. With the advancement of LLMs and RLLMs, the factors influencing their financial question-answering capabilities remain under-explored. Inspired by this gap, we systematically investigate the following issues: ($i$) the impact of prompting methods on LLMs/RLLMs; ($ii$) the influence of agentic frameworks on LLMs/RLLMs; and (iiiiii) whether multilingual alignment methods can enhance the multilingual financial question-answering abilities of LLMs/RLLMs. We utilize five LLMs and four RLLMs, conducting detailed experiments across seven representative methods.
@@ 3.1 Exploration of prompting methods
Current research extensively demonstrates that different prompting methods can significantly alter the performance of LLMs (wei2022chain; wang2023plan; zhang-etal-2024-wrong). The prompting method modifies certain aspects of the thinking prompt while ensuring the question remains intact, enabling LLMs to solve the question by following the structured steps of the thinking prompt.
We select three representative prompting methods to investigate the impact of varying prompt word changes on the performance of LLMs and RLLMs: Direct, Zero-shot CoT (kojima2022large), and Plan-and-Solve (wang2023plan). These prompting methods can be defined as:
@@ 3.2 Exploration of agentic frameworks
To further enhance the performance of LLMs, researchers have started to incorporate agentic systems. Specifically, by enabling LLMs to assume different roles and engage in continuous communication and cooperation, their performance can be significantly improved (luo2025large).
We select two agentic frameworks, and aim to explore the impact of LLMs and RLLMs in different agent collaborations: Self-Refine (madaan2023self) and S3 Agent (10.1145/3690642). These agentic frameworks can be defined as:
@@ 3.3 Exploration of multilingual alignment methods
We evaluate the performance of four standard multilingual approaches, including state-of-the-art cross-lingual prompting methods, on multilingual alignment: Direct, Translate-en (shi2023language), and Cross-lingual Prompting (qin-etal-2023-cross).
@@
  • Translate-en: Translate-en employs translation to convert the local language question and context into English, thereby preventing performance degradation of LLMs due to insufficient multilingual capabilities.
@@
Model Method Arithmetic Non-Arithmetic Overall Overall Easy Medium Hard Overall Easy Medium Hard Llama-3.1-8B-Instruct Direct grattafiori2024llama 10.88 12.66 9.89 9.28 23.44 22.45 32.10 20.27 16.50 (grattafiori2024llama) Zero-shot CoT kojima2022large 13.20 16.15 10.95 11.07 31.72 32.62 34.73 29.95 21.49 Plan-and-Solve wang2023plan 16.09 18.99 14.84 13.17 29.77 32.62 30.52 27.92 22.21 Self-Refine madaan2023self 17.95 22.70 13.07 15.56 32.06 34.74 35.26 29.27 24.26 S3 Agent 10.1145/3690642 19.62 22.27 15.54 19.46 30.80 37.28 32.63 26.57 24.62 GPT-4o-mini Direct hurst2024gpt 32.46 41.70 30.38 21.55 42.52 50.42 50.52 34.90 36.96 (hurst2024gpt) Zero-shot CoT kojima2022large 34.13 41.04 34.27 24.55 45.17 56.35 52.10 36.26 39.07 Plan-and-Solve wang2023plan 36.18 44.97 33.92 26.04 45.74 56.77 51.05 37.61 40.46 Self-Refine madaan2023self 33.30 41.26 33.21 22.45 42.06 52.11 45.78 35.13 37.22 S3 Agent 10.1145/3690642 30.88 38.42 28.26 22.75 41.72 51.69 46.84 34.23 35.73 Gemini-1.5-flash Direct team2024gemini 37.20 47.16 35.33 25.14 45.28 58.05 49.47 36.71 40.82 (team2024gemini) Zero-shot CoT kojima2022large 37.39 46.72 36.04 25.74 46.09 58.89 51.57 36.93 41.28 Plan-and-Solve wang2023plan 37.76 46.50 39.57 24.25 47.12 55.50 55.26 39.18 41.95 Self-Refine madaan2023self 34.41 42.35 33.92 23.95 42.52 52.96 50.00 33.78 38.04 S3 Agent 10.1145/3690642 37.20 45.85 34.62 27.54 44.59 57.20 50.52 35.36 40.51 Qwen-2.5-32B Direct yang2024qwen2 41.86 50.65 40.98 30.53 48.62 59.32 56.31 39.63 44.88 (yang2024qwen2) Zero-shot CoT kojima2022large 42.32 50.21 40.98 32.63 50.80 59.32 60.00 42.34 46.11 Plan-and-Solve wang2023plan 39.44 44.97 43.46 28.44 49.77 60.59 55.78 41.44 44.06 Self-Refine madaan2023self 41.58 49.56 42.75 29.64 49.65 57.62 54.73 43.24 45.19 S3 Agent 10.1145/3690642 42.51 51.52 41.34 31.13 48.85 61.44 56.31 38.96 45.34 DeepSeek-V3 Direct liu2024deepseek 56.27 65.50 58.65 41.61 62.06 73.72 69.47 52.70 58.86 (liu2024deepseek) Zero-shot CoT kojima2022large 56.18 63.97 59.71 42.51 61.83 71.18 70.52 53.15 58.71 Plan-and-Solve wang2023plan 55.16 63.97 56.18 42.21 63.33 70.76 71.05 56.08 58.81 Self-Refine madaan2023self 53.86 61.79 55.83 41.31 59.54 69.06 64.73 52.25 56.40 S3 Agent 10.1145/3690642 54.69 61.57 56.89 43.41 59.42 71.18 66.31 50.22 56.81 DeepSeek-R1-Distill-Qwen-32B Direct guo2025deepseek 50.41 60.48 50.17 36.82 57.12 65.25 64.73 49.54 53.41 (guo2025deepseek) Zero-shot CoT kojima2022large 51.25 60.91 51.94 37.42 56.55 68.64 63.15 47.29 53.62 Plan-and-Solve wang2023plan 51.16 60.91 50.53 38.32 55.86 66.52 62.63 47.29 53.26 Self-Refine madaan2023self 44.65 52.18 47.70 31.73 52.06 57.20 60.00 45.94 47.96 S3 Agent 10.1145/3690642 51.25 59.60 51.94 39.22 58.04 69.49 65.26 48.87 54.29 GPT-OSS-20B Direct agarwal2025gpt 50.04 61.57 48.40 35.62 57.93 60.59 67.89 52.25 53.57 (agarwal2025gpt) Zero-shot CoT kojima2022large 50.04 60.69 51.59 34.13 56.89 64.40 68.94 47.74 53.11 Plan-and-Solve wang2023plan 49.67 59.82 49.46 35.92 55.28 63.98 59.47 48.87 52.18 Self-Refine madaan2023self 48.27 56.55 48.76 36.52 57.58 69.49 61.57 49.54 52.44 S3 Agent 10.1145/3690642 52.00 61.79 51.23 39.22 58.62 68.22 63.68 51.35 54.96 Qwen-3-14B Direct yang2025qwen3 54.88 66.59 58.65 35.62 59.08 72.03 70.52 47.29 56.76 (yang2025qwen3) Zero-shot CoT kojima2022large 55.72 68.34 56.18 38.02 59.19 69.91 70.52 48.64 57.27 Plan-and-Solve wang2023plan 54.97 66.81 58.65 35.62 57.81 71.18 68.94 45.94 56.24 Self-Refine madaan2023self 54.41 65.93 56.89 36.52 58.27 71.61 66.84 47.52 56.14 S3 Agent 10.1145/3690642 54.13 65.06 57.24 36.52 58.50 71.18 67.36 47.97 56.09 O4-mini Direct o4-mini 61.30 70.08 65.01 46.10 70.22 77.54 75.78 63.96 65.29 (o4-mini) Zero-shot CoT kojima2022large 62.60 69.86 64.66 50.89 71.37 77.11 77.89 65.54 66.52 Plan-and-Solve wang2023plan 61.67 69.86 64.31 48.20 72.41 77.96 77.89 67.11 66.47 Self-Refine madaan2023self 61.20 69.21 60.77 50.59 70.34 75.00 74.73 65.99 65.29 S3 Agent 10.1145/3690642 61.76 69.65 62.19 50.59 71.49 76.69 75.78 66.89 66.11
@@ 4.1 Experimental Settings
We use the standard financial question answering benchmark FAMMA (xue2024famma) for experiments. FAMMA is a benchmark for financial multilingual multi-modal question answering, which includes English, Chinese, and French. Since current RLLMs are unimodal and cannot process multimodal information, we utilize the Basic Txt dataset in FAMMA, which converts multimodal data into textual information using OCR. The final dataset used for evaluation comprises 1,945 entries.
We conduct experiments on nine backbones, including five LLMs and four RLLMs: Meta-Llama-3.1-8B-Instruct (grattafiori2024llama), GPT-4o-mini (hurst2024gpt), Gemini-1.5-flash (team2024gemini), Qwen-2.5-32B (yang2024qwen2), DeepSeek-V3 (liu2024deepseek), DeepSeek-R1-Distill-Qwen-32B (guo2025deepseek), GPT-OSS-20B (agarwal2025gpt), think mode of Qwen-3-14B (yang2025qwen3), O4-mini (o4-mini). All open questions are evaluated using GPT-4o-mini based on correct answers. Due to computational constraints, all models were evaluated using a single run.
@@ 4.2 Analysis of Prompting Methods
Larger LLMs exhibit increased robustness to various prompting methods. It is evident that for larger LLMs, such as Gemini-1.5-flash and DeepSeek-V3, the performance gap between different prompting methods is not significant, with performance fluctuations remaining under 2%. This indicates that LLMs with superior performance are less affected by input prompts, demonstrating greater robustness to variations in input.
Performance Direct Translate-en (shi2023language) Self-Refine (madaan2023self) S3 Agent (10.1145/3690642) CLP (qin-etal-2023-cross) Meta-Llama-3.1-8B-Instruct (grattafiori2024llama) Chinese 16.54 16.54 (+0.00) 18.97 (+2.43) 19.22 (+2.68) 23.35 (+6.81) French 15.38 16.66 (+1.28) 13.46 (-1.92) 21.15 (+5.77) 17.30 (+1.92) Overall (Chinese+French) 16.22 16.57 (+0.35) 17.46 (+1.24) 19.75 (+3.53) 21.69 (+5.47) GPT-4o-mini (hurst2024gpt) Chinese 33.09 34.54 (+1.45) 34.79 (+1.70) 31.63 (-1.46) 38.19 (+5.10) French 28.84 30.76 (+1.92) 28.84 (+0.00) 32.69 (+3.85) 32.05 (+3.21) Overall (Chinese+French) 31.92 33.50 (+1.58) 33.15 (+1.23) 31.92 (+0.00) 36.50 (+4.58) Gemini-1.5-flash (team2024gemini) Chinese 40.38 38.68 (-1.70) 36.73 (-3.65) 39.65 (-0.73) 37.22 (-3.16) French 27.56 29.48 (+1.92) 26.92 (-0.64) 27.56 (+0.00) 29.48 (+1.92) Overall (Chinese+French) 36.86 36.15 (-0.71) 34.03 (-2.83) 36.33 (-0.53) 35.09 (-1.77) Qwen-2.5-32B (yang2024qwen2) Chinese 41.84 44.52 (+2.68) 44.28 (+2.44) 42.82 (+0.98) 43.30 (+1.46) French 29.48 29.48 (+0.00) 33.97 (+4.49) 35.89 (+6.41) 35.25 (+5.77) Overall (Chinese+French) 38.44 40.38 (+1.94) 41.44 (+3.00) 40.91 (+2.47) 41.09 (+2.65) DeepSeek-V3 (liu2024deepseek) Chinese 57.42 58.63 (+1.21) 55.96 (-1.46) 57.66 (+0.24) 61.31 (+3.89) French 50.00 50.00 (+0.00) 47.43 (-2.57) 44.23 (-5.77) 57.69 (+7.69) Overall (Chinese+French) 55.37 56.26 (+0.89) 53.61 (-1.76) 53.96 (-1.41) 60.31 (+4.94) DeepSeek-R1-Distill-Qwen-32B (guo2025deepseek) Chinese 48.41 49.87 (+1.46) 46.47 (-1.94) 52.06 (+3.25) 50.85 (+2.44) French 49.35 42.94 (-6.41) 44.87 (-1.16) 49.35 (+0.00) 52.56 (+3.21) Overall (Chinese+French) 48.67 47.97 (-0.70) 46.03 (-2.64) 51.32 (+2.65) 51.32 (+2.65) GPT-OSS-20B (agarwal2025gpt) Chinese 50.36 51.33 (+0.97) 50.60 (+0.24) 54.50 (+4.14) 54.76 (+4.40) French 51.28 46.15 (-5.13) 46.15 (-5.13) 55.76 (+4.48) 45.51 (-5.77) Overall (Chinese+French) 50.61 49.91 (-0.70) 49.38 (-1.23) 54.85 (+4.24) 52.21 (+1.60) Qwen-3-14B (yang2025qwen3) Chinese 57.66 56.69 (-0.97) 55.23 (-2.43) 55.47 (-2.19) 53.77 (-3.89) French 48.71 50.00 (+1.29) 42.3 (-6.41) 48.71 (+0.00) 51.28 (+2.57) Overall (Chinese+French) 55.20 54.85 (-0.35) 51.67 (-3.53) 53.61 (-1.59) 53.08 (-2.12) O4-mini (o4-mini) Chinese 65.45 65.93 (+0.48) 64.23 (-1.22) 67.63 (+2.18) 64.54 (-0.91) French 62.17 61.53 (-0.64) 60.25 (-1.92) 61.53 (-0.64) 63.46 (+1.29) Overall (Chinese+French) 64.55 64.72 (+0.17) 63.13 (-1.42) 65.69 (+1.14) 64.24 (-0.31)
@@
For RLLMs, competitive results can be achieved without the use of prompting methods. Benefiting from the Long CoT process, RLLMs can generate complex reasoning independently, without deliberate guidance from prompting methods. This renders prompting, which is effective for LLMs, less effective for RLLMs and may even diminish their performance. As shown in Table 1, Plan-and-Solve demonstrates poor performance on the four RLLMs.